OpenStudy (anonymous):

Just check my answers please? <3

OpenStudy (anonymous):

Write the point slope form of an equation for a line that passes through each point with the given slope. 15. (3,8) m=2 My answer: y-8=2x-6 18.(-6,1) m=-4 My answer: y-1=-4x+24 21. (8,-3) m=3/4 My answer: y-(-3)=3x/4-8 24. (9,-5) m=0 My answer: y-5=0x-0

OpenStudy (is3535):

good job in all

OpenStudy (anonymous):

They are all correct??

OpenStudy (is3535):

yea good job

OpenStudy (anonymous):

@kidrah69 help pleasse?

OpenStudy (loser66):

15) it is correct but not the slope form which is y = mx +b, hence, you need simplify more, the answer should be y = 2x +2

OpenStudy (anonymous):

Thanks(:

OpenStudy (loser66):

18) Same as 15) but check the sign of 24, the final answer should be y = -4x -23

OpenStudy (loser66):

Let me show you the steps for 21) m = (3/4) , point ( 8, -3) our formula is y = mx +b hence b = y - mx replace all, b = -3 - (3/4)*8 = -3 -6 = -9 now, plug back to the formula y = (3/4) x -9

OpenStudy (loser66):

the last one, if m =0, that is the line is horizontal line. We just take the value of y y = -5

OpenStudy (anonymous):

Thankssssss